# Efficient classical algorithms for linear optical circuits

### v2

@@ Abstract
We present efficient classical algorithms to approximate expectation values and probability amplitudes in linear optical circuits. Specifically, our algorithm efficiently approximates the expectation values of observables in linear optical circuits for arbitrary product input states within an additive error under a mild condition. This result suggests that certain applications of linear optical circuits relying on expectation value estimation, such as photonic variational algorithms, may face challenges in achieving quantum advantage. In addition, the (marginal) output probabilities of boson sampling with arbitrary product input states can be efficiently approximated using our algorithm, implying that boson sampling can be efficiently simulated if its output probability distribution is polynomially sparse. Our method generalizes Gurvits’s algorithm, originally designed to approximate the permanent, to also approximate the hafnian of complex symmetric matrices with an additive error. The algorithm also solves a molecular vibronic spectra problem for arbitrary product input states as precisely as boson samplers. Finally, our method extends to near-Clifford circuits, enabling the classical approximation of their expectation values of any observables and (marginal) output probabilities.
@@
Introduction.— Linear optical system has been considered as a promising platform for universal quantum computing [1] and also for demonstrating the quantum computational advantage in the near future since the seminal work on boson sampling [2]. The boson sampling proposal [2, 3, 4] motivated various experiments to demonstrate the quantum advantage [5, 6, 7, 8, 9] while numerous classical algorithms have also been developed to challenge them, leading to significant advancements in understanding the computational complexity of linear optical circuits [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22].
Lately, the main direction of this research has been shifting toward achieving practical quantum advantage: finding a useful problem that cannot be efficiently solved by classical computers but can be addressed by exploiting linear optical circuits [23, 24, 25, 26, 27, 28, 29, 30, 31]. One promising direction of exploiting sampling devices involves problems where the goal is to compute physical quantities through sampling, such as in Monte Carlo methods. Especially, many such problems can be formulated as estimating expectation values of observables, often referred to as the quantum mean value problem (e.g., Ref. [32]). Although various proposals have been suggested in this direction with boson sampling [23, 28, 29, 30, 31], no provable advantage has yet been demonstrated; thus, these approaches cannot rule out the existence of efficient classical algorithms. For example, the molecular vibronic spectra problem, which has been considered a promising potential application, can be efficiently solved using Gaussian boson sampling [23, 3, 4], and several experiments have been conducted to validate this proposal [33, 34, 35]. However, a recent result shows that a classical algorithm can achieve comparable performance [36], ruling out the practical advantage of boson sampling for this problem.
The biggest challenge in leveraging the quantum advantage of boson sampling for solving the associated quantum mean value problem is the limited accuracy due to the finite sample size, as in the above example. This limitation ultimately leaves room for a classical algorithm to achieve the same performance. Thus, identifying the classical limit on the achievable accuracy for the quantum mean value problem is crucial to rigorously establish a practical quantum advantage. Although classical algorithms for the quantum mean value problem have been developed in various contexts [32, 37, 38, 39, 40, 41], no such algorithm has yet been studied for linear optical circuits, to the best of our knowledge.
@@
Figure 1: Summary of classical algorithms introduced in the main text. (a) A schematic of the general expectation value of an $M$-mode linear optical circuit (Theorem 1) or an $M$-qubit near-Clifford circuit (Theorem 3). If $\hat{O}_{i}$ are photon-number projectors $|m_{i}\rangle\langle m_{i}|_{i}$, the expectation values represent marginal output probabilities of a boson sampling circuit, resulting in an efficient simulation of a sparse boson sampling (Corollary 1). (b) A probability amplitude of a linear optical circuit stands for generalized Gurvits’s algorithm (Theorem 2). When $|\phi_{i}\rangle$’s are single-photon Fock states and $|\psi_{i}\rangle$’s are squeezed vacuum states, the amplitude is proportional to the hafnian of a general complex symmetric matrix (Corollary 2). A probability amplitude is reduced to a phase shifter expectation value, Eq. (8), with $|\phi_{i}\rangle=|\psi_{i}\rangle$ and $\hat{U}\rightarrow\hat{U}^{\dagger}e^{i\hat{\bm{n}}\cdot\bm{\phi}}\hat{U}$ (Corollary 3).
In this work, we present an efficient classical algorithm for approximating the expectation values in linear optical circuits. Our algorithm works for arbitrary product input states when the norm of the operator of interest is well-bounded, dequantizing a wide range of potential applications, such as computing cost functions in variational quantum algorithms [42, 43]. In addition, the algorithm enables an efficient approximation of the (marginal) output probabilities of boson sampling circuits for arbitrary input states. Thus, it proves that any boson sampling can be classically simulated if the output probability distribution is polynomially sparse [44], extending previous results that applied only to restricted cases [45, 46, 47].
We further extend our result by identifying cases in which it remains valid even for operators with unbounded norm. One such case is the phase operator in the generalized molecular vibronic spectra problem, which was proposed as a candidate achieving practical quantum advantage [36]. We fully solve this problem by providing a classically efficient algorithm for arbitrary input states. Additionally, it can be used for approximating binned probabilities of linear optical circuits with arbitrary input states, which have been proposed for verifying boson sampling experiments [48, 49, 50, 51]. In the language of computational complexity, our work generalizes Gurvits’s algorithm [52, 2]. It provides a quantum-inspired classical algorithm to approximate matrix functions, such as hafnian [53], by mapping the matrix function of interest to a Gaussian boson sampling’s output amplitude [3, 54].
@@
Finally, using a similar procedure, we apply our method to near-Clifford circuits [55, 56] and show that we can efficiently estimate the expectation values of observables and any (marginal) output probabilities for these circuits within an additive error.
@@
Efficient approximation of expectation values.— Let $|\psi\rangle=\otimes_{i=1}^{M}|\psi_{i}\rangle$ be a product input state in an $M$-mode(-qubit) system, and $\hat{U}$ be a quantum circuit that is applied to the input state. For the output state of the circuit, $\hat{U}|\psi\rangle$, we consider the expectation value of a product operator $\hat{O}=\otimes_{i=1}^{M}\hat{O}_{i}$ (see Fig. 1(a)):
@@
Throughout this work, input states $|\psi\rangle$ and operators $\hat{O}$ are assumed to be in a product form unless stated otherwise, which is practically relevant in various applications. When an operator acts nontrivially only on a subsystem $A$ and trivially on the rest of the system $B$, we write the operator as $\hat{O}=\hat{O}_{A}\otimes\hat{\mathbb{1}}_{B}=\bigotimes_{i\in A}\hat{O}_{i}\otimes\bigotimes_{i\in B}\hat{\mathbb{1}}_{i}$ and define the reduced density matrix $\hat{\rho}_{A}\equiv\mbox{Tr}_{B}[\hat{U}|\psi\rangle\langle\psi|\hat{U}^{\dagger}]$.
We first focus on a linear optical circuit $\hat{U}$. Our first main result is a classical algorithm that efficiently approximates the expectation values of operators in a linear optical circuit:
Theorem 1 (Expectation value approximation in linear optical circuit).
Consider an $M$-mode linear optical circuit $\hat{U}$ and an operator $\hat{O}$. The expectation values $\langle\psi|\hat{U}^{\dagger}\hat{O}\hat{U}|\psi\rangle$ can be approximated within additive error $\epsilon$ with probability $1-\delta$ in running time $O(M^{2}\lVert\hat{O}_{A}\rVert^{2}_{2}\|\hat{\rho}_{A}\|_{2}^{2}\log(1/\delta)/\epsilon^{2})$.
@@
$\langle\psi|\hat{U}^{\dagger}\hat{O}\hat{U}|\psi\rangle$ $=\frac{1}{\pi^{l}}\int d^{2l}\alpha\chi_{\hat{O}_{A}}(\bm{\alpha})\chi_{\hat{\rho}_{A}}^{*}(\bm{\alpha})$ (2)
$=\int d^{2l}\alpha\frac{|\chi_{\hat{O}_{A}}(\bm{\alpha})|^{2}}{\pi^{l}\|\hat{O}_{A}\|_{2}^{2}}\frac{\|\hat{O}_{A}\|_{2}^{2}\chi_{\hat{\rho}_{A}}^{*}(\bm{\alpha})}{\chi^{*}_{\hat{O}_{A}}(\bm{\alpha})},$ (3)
@@
We then show that the median-of-means estimator [64] yields an estimate with additive error $\epsilon$ with probability $1-\delta$ when $O(\|\hat{O}_{A}\|_{2}^{2}\|\hat{\rho}_{A}\|_{2}^{2}\log\delta^{-1}/\epsilon^{2})$ number of samples is given, by showing that the variance of the random variable $X(\bm{\alpha})$ is upper-bounded by $\|\hat{O}_{A}\|_{2}^{2}\|\hat{\rho}_{A}\|_{2}^{2}$. Note that the median-of-means estimator for complex random variables can be constructed by combining the real and imaginary parts estimators [57]. Since it takes $O(M^{2})$ time for computing random variable $X(\bm{\alpha})$, the total running time is $O(M^{2}\lVert\hat{O}_{A}\rVert^{2}_{2}\|\hat{\rho}_{A}\|_{2}^{2}\log(1/\delta)/\epsilon^{2})$. ∎
Therefore, we achieve an efficient additive estimation of the expectation value in polynomial time when $\lVert\hat{O}\rVert^{2}_{2}=O(\text{poly}(M))$, a condition that holds in many interesting cases, as elaborated below.
Theorem 1 has various intriguing implications. First, any applications of linear optical circuits, such as boson sampling, that rely on estimating a physical observable $\hat{O}$ can be dequantized if the target observable’s norm is properly bounded; thus, this establishes a boundary for a classically simulable regime when it comes to approximating expectation values. Furthermore, when $\hat{O}$ is an orthogonal projector, its expectation value directly corresponds to the associated output probability. Therefore, if the rank of the orthogonal projector is well-bounded, the output probability of a linear optical circuit can be efficiently approximated. A notable example is a photon-number projector in boson sampling [2]. In this case, Theorem 1 implies that the (marginal) output probabilities of boson sampling with arbitrary input states can be efficiently approximated. To the best of our knowledge, such a result has only been known for the Fock state input [2, 49, 36]. Remarkably, this finding also enables an efficient simulation of boson sampling with arbitrary input states if polynomially many outcomes dominate the output probabilities [65, 44, 66].
@@
Boson sampling with an arbitrary input state can be efficiently simulated if its output probability distribution is polynomially sparse.
In addition, we note that quantum sampling with a peaked output probability distribution, i.e., one that has an outcome with an inverse-polynomially large probability, has drawn attention for a verifiable quantum advantage of sampling [67, 68] because the peak can serve as a verification. However, our result, together with the algorithm in Refs. [65, 44], indicates that the peak of the output probability distribution in boson sampling with any input states can be efficiently identified using classical computers. We emphasize that Corollary 1 is valid for a general case regardless of phase space negativity [69, 47] or stellar rank [70] of the input state, resources for quantum advantages.
Another interesting example is recently proposed photonic variational quantum algorithms [42, 43, 71], whose cost function is written in terms of the expectation value as
$f(\theta)=\langle\psi|\hat{V}_{2}^{\dagger}e^{-i\theta\hat{n}_{k}}\hat{V}_{1}^{\dagger}\hat{O}\hat{V}_{1}e^{i\theta\hat{n}_{k}}\hat{V}_{2}|\psi\rangle,$ (4)
where $\hat{V}_{1}$ and $\hat{V}_{2}$ are linear optical circuits and $\hat{n}_{k}$ is the number operator on the $k$th mode. The main subroutine for obtaining the gradient is the estimation of $f(\theta)$’s for several values of $\theta$ by running the corresponding linear optical circuits. Since $\hat{V}_{1}e^{i\theta\hat{n}_{k}}\hat{V}_{2}$ is also a linear optical circuit, the estimation of the cost function can be replaced by the classical algorithm of Theorem 1 with $\hat{U}=\hat{V}_{1}e^{i\theta\hat{n}_{k}}\hat{V}_{2}$. Thus, our classical algorithm can replace the quantum algorithms when the norm of the operator $\hat{O}$ is in a product form and bounded.
Generalized Gurvits’s algorithm.— We now present another classical algorithm approximating a probability amplitude $\langle\phi|\hat{U}|\psi\rangle$ (See Fig. 1 (b)). Note that when $|\phi\rangle$ and $|\psi\rangle$ are single-photon Fock states, the amplitude becomes a permanent [2], which can be efficiently approximated by Gurvits’s algorithm [52]. Hence, we call our classical algorithm a generalized Gurvits’s algorithm, which fully extends the original Gurvits’s algorithm in the context of quantum optics:
Theorem 2 (Generalized Gurvits’s algorithm).
The probability amplitude $\langle\phi|\hat{U}|\psi\rangle$ of a linear optical circuit $\hat{U}$ with arbitrary states $|\psi\rangle$ and $|\phi\rangle$ can be approximated within additive error $\epsilon$ with probability $1-\delta$ in running time $O(M^{2}\log\delta^{-1}/\epsilon^{2})$.
@@
$\langle\phi|\hat{U}|\psi\rangle$ $=\frac{1}{\pi^{M}}\int d^{2M}\alpha\langle\phi|\hat{U}|\bm{\alpha}\rangle\langle\bm{\alpha}|\psi\rangle$ (5)
$=\int d^{2M}\alpha\frac{\langle\phi|\hat{U}|\bm{\alpha}\rangle}{\langle\bm{\alpha}|\psi\rangle^{*}}\frac{|\langle\bm{\alpha}|\psi\rangle|^{2}}{\pi^{M}},$ (6)
where $|\bm{\alpha}\rangle$ represents a coherent state. Thus, if we sample $\bm{\alpha}$ from $|\langle\bm{\alpha}|\psi\rangle|^{2}/\pi^{M}$, which is a proper probability distribution (Husimi Q-distribution of the state $|\psi\rangle$ [72]) and easy to sample from, and average $\langle\phi|\hat{U}|\bm{\alpha}\rangle/\langle\bm{\alpha}|\psi\rangle^{*}$ over the samples, we obtain an estimate of the amplitude. We prove that this procedure is efficient and gives an estimate with $\epsilon$ additive error with probability $1-\delta$ in running time $O(M^{2}\log\delta^{-1}/\epsilon^{2})$ in SM Sec. S3 [57]. ∎
Note that the original Gurvits’s algorithm applies only when $|\psi\rangle=|\bm{m}\rangle$, $|\phi\rangle=|\bm{n}\rangle$ with $\bm{n},\bm{m}\in\{0,1\}^{M}$, and its variants [73, 36] apply when $\bm{m}\in\{0,1\}^{M}$ or $\bm{n}=\bm{m}$, corresponding to approximating the permanent of matrices with repeated rows and columns in a certain pattern. Not only does our algorithm extend to computing the permanent of matrices with repeated rows and columns in a general pattern, but it also introduces a new classical approach to approximating the permanent using a quantum-optics-inspired method.
Our algorithm can also approximate the hafnian [53], a generalization of the permanent, by using the fact that the hafnian of any $M\times M$ complex symmetric matrix $R$ ($M$ is even) with the largest singular value $\lambda_{\text{max}}<1$ can be expressed as a probability amplitude of an $M$-mode Gaussian boson sampling circuit [3, 54]:
@@
where $|\bm{r}\rangle=\otimes_{i=1}^{M}|r_{i}\rangle$ denotes a product of squeezed vacuum states $|r_{i}\rangle$ of squeezing parameter $r_{i}$, $|\bm{1}\rangle=\otimes_{i=1}^{M}|1\rangle$ is $M$ single-photon states, and $Z\equiv\prod_{i=1}^{M}\cosh{r_{i}}=\prod_{i=1}^{M}(1-\lambda_{i}^{2})^{-\frac{1}{2}}$ with $R$’s singular values $\lambda_{i}$. Until now, efficient additive approximation algorithms have been known only for restricted cases [47, 36]. Notably, our generalized Gurvits’s algorithm leads to an additive approximation algorithm for the hafnian without any constraints, which can be seen by substituting $|\psi\rangle=|\bm{r}\rangle$, $|\phi\rangle=|\bm{1}\rangle$ in Theorem 2:
@@
The hafnian of an $M\times M$ complex symmetric matrix $R$ can be approximated within additive error $\epsilon\norm{R}^{M/2}$ with probability $1-\delta$ in $O(M^{3}\log\delta^{-1}/\epsilon^{2})$ time.
A detailed proof is provided in SM Sec. S4 [57]. By a similar method, the loop hafnian of an arbitrary complex symmetric matrix [54] also can be efficiently approximated.
It is worth emphasizing that Theorem 2 can also be used to approximate the overlap of two quantum states generated by linear optical circuits. Such a quantity has attracted attention for machine learning applications of boson sampling, such as kernel method [29, 30, 31]. More specifically, boson samplers are used to estimate the overlap of two quantum states corresponding to two different data vectors $\bm{x},\bm{y}$ as $|\langle\psi|\hat{U}^{\dagger}(\bm{x})\hat{U}(\bm{y})|\psi\rangle|^{2}$, where $\hat{U}(\bm{x})$ corresponds to the linear optical circuits whose configuration depend on the data vector $\bm{x}$. Thus, our classical method can replace boson samplers for this purpose.
Phase shifter.— We now consider a problem of approximating the expectation value of a phase shifter $e^{i\hat{\bm{n}}\cdot\bm{\phi}}$ of a linear optical circuit [74]
@@
where $\hat{\bm{n}}\equiv(\hat{n}_{1},\dots,\hat{n}_{M})$ and $\bm{\phi}\in\mathbb{R}^{M}$ represents the phase vector. The expectation value of a phase shifter has various applications such as the molecular vibronic spectra problem [23, 36] and approximating binned probability, which has been proposed for verifying boson sampling experiments [48, 49, 50, 51].
In particular, the molecular vibronic spectra problem [23] has been considered as a potential application of boson sampling, which is essentially equivalent to computing grouped output probabilities $p(\bm{m})$ of boson sampling (See SM Sec. S4 [57] for more details):
@@
where $\bm{\omega}\in\mathbb{Z}^{M}_{\geq 0}$ and $\Omega=\{0,\dots,\Omega_{\text{max}}\}$. While a boson sampler can efficiently approximate Eq. (9) by sampling and counting the frequency of each group indexed by $\Omega$, classical approaches seem unlikely to be promising not just because the output probability $p(\bm{m})$ is #P-hard to compute but also because the number of outcomes $\bm{m}$ corresponding to each group $\Omega$ is generally exponentially large. In Ref. [36], however, the authors observe that the Fourier components of Eq. (9) are expressed as the expectation value of a phase shifter as Eq. (8). Therefore, if Eq. (8) is additively approximated, the grouped probabilities Eq. (9) are reproduced within an additive error through the inverse Fourier transform [57].

### v3

# Classical algorithms for estimating expectation values in linear-optical circuits

@@ Abstract
We present a classical algorithm for approximating the expectation values of observables in linear-optical circuits with arbitrary product input states, achieving additive-error accuracy. This result indicates that current applications of photonic systems aimed at demonstrating practical quantum supremacy through expectation value estimation, such as photonic variational algorithms, may face challenges in attaining the computational advantage. It also implies the output probabilities of boson sampling with arbitrary product input states can be efficiently approximated by our method, resulting that boson sampling becomes efficiently simulable when its output probability distribution is polynomially sparse. We also develop an efficient classical algorithm for estimating transition amplitudes of arbitrary product states in linear-optical circuits. This provides additive-error approximation algorithms for matrix functions associated with linear-optical circuits, such as the (loop-)hafnian, which are of independent interest. As an application, it solves the generalized molecular vibronic spectra problem (Oh et al., 2024), previously suggested as a candidate for practical quantum advantage. Finally, we extend our framework to near-Clifford circuits, enabling classical approximation of their expectation values.
@@
Introduction.— Linear-optical system has been considered as a promising platform for universal quantum computing [1] and also for demonstrating the quantum computational advantage in the near future since the seminal work on boson sampling [2]. The boson sampling proposal [2, 3, 4] motivated various experiments to demonstrate the quantum advantage [5, 6, 7, 8, 9] while numerous classical algorithms have also been developed to challenge them, leading to significant advancements in understanding the computational complexity of linear-optical circuits [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
Recently, the focus has shifted from demonstrating hardness through sampling to identifying practically motivated tasks that reflect a practical form of quantum advantage. A common viewpoint is that near-term quantum devices aim to estimate expectation values of observables with controlled additive accuracy, a task believed to be classically hard [24, 25, 26, 27, 28, 29, 30]. Within linear-optics, several proposals adopt this estimation-based framework [31, 32, 33, 34, 35], yet no provable advantage has been established. A notable case is molecular vibronic spectra (MVS): while Gaussian boson sampling enables an efficient photonic procedure [31, 3, 4, 36, 37, 38], recent work shows that a classical algorithm achieves comparable performance [39], because finite sampling enforces an additive-error tolerance, ruling out an advantage for MVS in this setting.
The essential issue, therefore, is precision: quantum devices estimate expectation values from finite samples, making additive-error approximations intrinsic to measurement (see SM Sec. S2 [40]). This immediately opens room for classical counterparts operating under the same additive tolerance. Accordingly, a rigorous assessment of practical quantum advantage requires identifying the classical limit of efficient estimability for the same observables within an additive-error tolerance. Although additive-error estimators for expectation values have been analyzed in other architectures [24, 41, 26, 42, 43, 44], to the best of our knowledge no such provable algorithm existed for linear-optical circuits.
@@
Figure 1: Summary of classical algorithms introduced in the main text. (a) A schematic of the general expectation value of an $M$-mode linear-optical circuit (Theorem 1). If $\hat{O}_{i}$ are photon-number projectors $|m_{i}\rangle\langle m_{i}|_{i}$, the expectation values represent marginal output probabilities of a boson sampling circuit, resulting in an efficient simulation of a sparse boson sampling (Corollary 1). (b) A transition amplitude of a linear-optical circuit (Theorem 2). When $|\phi_{i}\rangle$’s are single-photon Fock states and $|\psi_{i}\rangle$’s are squeezed vacuum states, the amplitude is proportional to the hafnian of a general complex symmetric matrix (Corollary 2). A transition amplitude is reduced to a phase shifter expectation value, Eq. (6), with $|\phi_{i}\rangle=|\psi_{i}\rangle$ and $\hat{U}\rightarrow\hat{U}^{\dagger}e^{i\hat{\bm{n}}\cdot\bm{\phi}}\hat{U}$ (Corollary 3).
In this work, we present a classical algorithm for approximating expectation values in linear-optical circuits within additive-error. When the observable $\hat{O}=\hat{O}_{A}\otimes\mathbb{1}_{B}$} has polynomially bounded Hilbert–Schmidt norm, $\|\hat{O}_{A}\|_{2}^{2}=O(\mathrm{poly}(M))$, the algorithm is efficient for arbitrary product input states. This dequantizes a broad class of estimation tasks relevant to NISQ photonic computing that align with the framework of practical quantum advantage, including photonic variational quantum algorithms (VQAs) where each term of the target Hamiltonian is a local observable [45, 46, 47, 48, 49, 50]. The same framework efficiently approximates (marginal) output probabilities of boson-sampling circuits with arbitrary product inputs. Hence, any boson sampling is classically simulable if its output distribution is polynomially sparse [51], generalizing earlier restricted results [52, 53, 54]. Consequently, in the polynomially sparse regime, peaks of inverse-polynomial probability can be classically identified, tempering peak-based proposals of verifiable quantum advantage based on linear optics [55, 56, 57, 51].
We further develop an efficient quantum-optics-inspired algorithm for transition amplitudes in linear-optical circuits with arbitrary product inputs. While the single-photon Fock case involves permanents, classically estimable via Gurvits’ algorithm [58, 2], our analysis follows a distinct route with comparable guarantees and extends to quantum-inspired estimators for other matrix functions, including the (loop-)hafnian [59, 3, 60]. It also provides a complete classical solution to the generalized MVS problem [39] and to efficient estimation of binned output probabilities used for boson-sampling verification [61, 62, 63, 64].
@@
Finally, we apply our method to near-Clifford circuits [65, 66] and show that we can efficiently estimate the expectation values of observables and any (marginal) output probabilities for these circuits within an additive-error.
@@
Approximation of expectation values.— Let $|\psi\rangle=\otimes_{i=1}^{M}|\psi_{i}\rangle$ be a product input state in an $M$-mode system, and $\hat{U}$ be a quantum circuit applied to the input state. For the output state of the circuit, $\hat{U}|\psi\rangle$, we consider the expectation value of a product operator $\hat{O}=\otimes_{i=1}^{M}\hat{O}_{i}$ (see Fig. 1(a)):
@@
Throughout this work, input states $|\psi\rangle$ and operators $\hat{O}$ are assumed in a product form unless stated otherwise, which is practically relevant in various applications. When an operator acts nontrivially only on a subsystem $A$ and trivially on the rest of the system $B$, we write the operator as $\hat{O}=\hat{O}_{A}\otimes\hat{\mathbb{1}}_{B}=\bigotimes_{i\in A}\hat{O}_{i}\otimes\bigotimes_{i\in B}\hat{\mathbb{1}}_{i}$ and define the reduced density matrix $\hat{\rho}_{A}\equiv\mbox{Tr}_{B}[\hat{U}|\psi\rangle\langle\psi|\hat{U}^{\dagger}]$.
We first focus on a linear-optical circuit $\hat{U}$. Our first main result is a classical algorithm that approximates the expectation values of operators in a linear-optical circuit:
Theorem 1 (Expectation-value approximation in linear-optical circuit).
Consider an $M$-mode linear-optical circuit $\hat{U}$ and an operator $\hat{O}$. The expectation values $\langle\psi|\hat{U}^{\dagger}\hat{O}\hat{U}|\psi\rangle$ can be approximated within additive-error $\epsilon$ with probability $1-\delta$ in running time $O(M^{2}\lVert\hat{O}_{A}\rVert^{2}_{2}\mbox{Tr}[\hat{\rho}_{A}^{2}]\log(1/\delta)/\epsilon^{2})$, where $\|\hat{O}_{A}\|_{2}^{2}=\mbox{Tr}[\hat{O}_{A}^{\dagger}\hat{O}_{A}]$.
@@
$\langle\psi|\hat{U}^{\dagger}\hat{O}\hat{U}|\psi\rangle$ $=\int d^{2l}\alpha\frac{|\chi_{\hat{O}_{A}}(\bm{\alpha})|^{2}}{\pi^{l}\|\hat{O}_{A}\|_{2}^{2}}\frac{\|\hat{O}_{A}\|_{2}^{2}\chi_{\hat{\rho}_{A}}^{*}(\bm{\alpha})}{\chi^{*}_{\hat{O}_{A}}(\bm{\alpha})},$ (2)
@@
We then show that the median-of-means estimator [73] yields an estimate with additive-error $\epsilon$ with probability $1-\delta$ when $O(\|\hat{O}_{A}\|_{2}^{2}\mbox{Tr}[\hat{\rho}_{A}^{2}]\log\delta^{-1}/\epsilon^{2})$ number of samples is given, by showing that the variance of the random variable $X(\bm{\alpha})$ is upper-bounded by $\|\hat{O}_{A}\|_{2}^{2}\mbox{Tr}[\hat{\rho}_{A}^{2}]$. Note that the median-of-means estimator for complex random variables can be constructed by combining the real and imaginary parts estimators [40]. Since it takes $O(M^{2})$ time for computing random variable $X(\bm{\alpha})$, the total running time is $O(M^{2}\lVert\hat{O}_{A}\rVert^{2}_{2}\mbox{Tr}[\hat{\rho}_{A}^{2}]\log(1/\delta)/\epsilon^{2})$. ∎
Therefore, we achieve an efficient additive estimation of the expectation value in polynomial time when $\lVert\hat{O}_{A}\rVert^{2}_{2}=O(\text{poly}(M))$. Such a condition holds whenever $\hat{O}_{A}$ has $O(\text{poly}(M))$ rank with polynomially large singular values. We present two important classes of observables that satisfy the condition.
A first important class arises when $\hat{O}_{A}$ is local observables, directly relevant to photonic VQAs implemented by linear-optical circuits [45, 46, 47, 48, 49, 74, 50]. In these schemes, the physically relevant target Hamiltonian decomposes into few-mode local terms,
$\hat{H}=\sum_{i}\alpha_{i}\hat{H}_{i}+\sum_{i<j}\beta_{ij}\hat{H}_{ij}+\cdots,$ (3)
and each $\hat{H}_{i},\hat{H}_{ij},\dots$ acts on a small number of modes with bounded Hilbert–Schmidt norm as $O(1)$. Our algorithm efficiently estimates the contribution of each local term, and the total energy follows by summation. The same bounded-norm observables also appear as the primary measurement targets in other NISQ paradigms, such as analog and digital quantum simulators [75, 76, 77, 78, 79, 80, 81], where physical quantities, such as magnetizations, correlations, or energy densities, are expressed as few-body expectation values. Hence, the condition $\|\hat{O}_{A}\|_{2}^{2}=O(\mathrm{poly}(M))$ precisely characterizes the physically relevant observables not only in photonic VQAs but broadly across NISQ devices, delineating a classically simulable regime for expectation-value estimation in linear optics (See SM Sec. S2 for more details [40].).
A second important class consists of observables where $\hat{O}_{A}$ is a rank-one projector of product form. This case is directly tied to projective measurements on the output of a linear-optical circuit. A notable example is a photon-number projector in boson sampling [2]. In this case, Theorem 1 implies that the (marginal) output probabilities of boson sampling with arbitrary input states can be efficiently approximated. To the best of our knowledge, such a result has only been known for the Fock state input [2, 62, 39]. Remarkably, this finding also enables an efficient simulation of boson sampling with arbitrary product input if polynomially many outcomes dominate the output probabilities [57, 51, 82]. (Note that sparsity of the output distribution can be determined. See SM Sec. S9 for more details [40].)
@@
Boson sampling with an arbitrary product input state can be efficiently simulated if its output probability distribution is polynomially sparse.
In addition, note that quantum sampling with a peaked output probability distribution, i.e., one that has an outcome with an inverse-polynomially large probability, has drawn attention for a verifiable quantum advantage of sampling [55, 56] because the peak can serve as a verification. However, our result, together with the algorithm in Refs. [57, 51], indicates that the peak of the output probability distribution in boson sampling with any product input can be efficiently identified using classical computers. We emphasize that Corollary 1 is valid for a general case regardless of phase space negativity [83, 54] or stellar rank [84] of the input state, resources for quantum advantages.
Approximation of transition amplitudes.— We now present another classical algorithm approximating a transition amplitude $\langle\phi|\hat{U}|\psi\rangle$ (See Fig. 1 (b)). Note that when $|\phi\rangle$ and $|\psi\rangle$ are single-photon Fock states, the amplitude becomes a permanent [2], which can be efficiently approximated by Gurvits’ algorithm [58]. In contrast, our algorithm works for arbitrary product inputs.
Theorem 2 (Transition amplitude approximation in linear-optical circuit).
The transition amplitude $\langle\phi|\hat{U}|\psi\rangle$ of a linear-optical circuit $\hat{U}$ with arbitrary states $|\psi\rangle$ and $|\phi\rangle$ can be approximated within additive-error $\epsilon$ with probability $1-\delta$ in running time $O(M^{2}\log\delta^{-1}/\epsilon^{2})$.
@@
$\langle\phi|\hat{U}|\psi\rangle$ $=\int d^{2M}\alpha\frac{\langle\phi|\hat{U}|\bm{\alpha}\rangle}{\langle\bm{\alpha}|\psi\rangle^{*}}\frac{|\langle\bm{\alpha}|\psi\rangle|^{2}}{\pi^{M}},$ (4)
where $|\bm{\alpha}\rangle$ represents a coherent state. Thus, if we sample $\bm{\alpha}$ from $|\langle\bm{\alpha}|\psi\rangle|^{2}/\pi^{M}$, which is a proper probability distribution (Husimi Q-distribution of the state $|\psi\rangle$ [85]) and easy to sample from, and average $\langle\phi|\hat{U}|\bm{\alpha}\rangle/\langle\bm{\alpha}|\psi\rangle^{*}$ over the samples, we obtain an estimate of the amplitude. We prove that this procedure is efficient and gives an estimate with $\epsilon$ additive-error with probability $1-\delta$ in running time $O(M^{2}\log\delta^{-1}/\epsilon^{2})$ in SM Sec. S4 [40]. ∎
Note that the original Gurvits’ algorithm applies only when $|\psi\rangle=|\bm{m}\rangle$, $|\phi\rangle=|\bm{n}\rangle$ with $\bm{n},\bm{m}\in\{0,1\}^{M}$, and its variants [86, 39] apply when $\bm{m}\in\{0,1\}^{M}$ or $\bm{n}=\bm{m}$, corresponding to approximating the permanent of matrices with repeated rows and columns in a certain pattern. Notably, our algorithm not only extends to computing the permanent of matrices with repeated rows and columns in a general pattern, but also introduces a new approach to approximating the permanent using a quantum-optics-inspired method.
Our algorithm can also approximate the hafnian [59], a generalization of the permanent, by using the fact that the hafnian of any $M\times M$ complex symmetric matrix $R$ ($M$ is even) with the largest singular value $\lambda_{\text{max}}<1$ can be expressed as a transition amplitude of an $M$-mode Gaussian boson sampling circuit [3, 60]:
@@
[…]
@@
The hafnian of an $M\times M$ complex symmetric matrix $R$ can be approximated within additive-error $\epsilon\norm{R}^{M/2}$ with probability $1-\delta$ in $O(M^{3}\log\delta^{-1}/\epsilon^{2})$ time.
A detailed proof is provided in SM Sec. S5 [40]. By a similar method, the loop hafnian of an arbitrary complex symmetric matrix [60] can also be efficiently approximated.
It is worth emphasizing that Theorem 2 can also be used to approximate the overlap of two quantum states generated by linear-optical circuits. Such a quantity has attracted attention for machine learning applications of boson sampling, such as kernel method [33, 34, 35]. More specifically, boson samplers are used to estimate the overlap of two quantum states corresponding to two different data vectors $\bm{x},\bm{y}$ as $|\langle\psi|\hat{U}^{\dagger}(\bm{x})\hat{U}(\bm{y})|\psi\rangle|^{2}$, where $\hat{U}(\bm{x})$ corresponds to the linear-optical circuits whose configuration depend on the data vector $\bm{x}$. Thus, our classical method can replace boson samplers for this purpose.
Phase shifter.— We now consider a problem of approximating the expectation value of a phase shifter $e^{i\hat{\bm{n}}\cdot\bm{\phi}}$ of a linear-optical circuit [87]
@@
where $\hat{\bm{n}}\equiv(\hat{n}_{1},\dots,\hat{n}_{M})$ and $\bm{\phi}\in\mathbb{R}^{M}$ represents the phase vector. The expectation value of a phase shifter has various applications such as the MVS problem [31, 39] and approximating binned probability, proposed for verifying boson sampling experiments [61, 62, 63, 64].
In particular, the MVS problem [31] has been considered as a potential application of boson sampling, which is essentially equivalent to computing grouped output probabilities $p(\bm{m})$ of boson sampling (See SM Sec. S4 [40] for more details):
@@
where $\bm{\omega}\in\mathbb{Z}^{M}_{\geq 0}$ and $\Omega=\{0,\dots,\Omega_{\text{max}}\}$. In Ref. [39], however, the authors observe that the Fourier components of Eq. (7) are expressed as the expectation value of a phase shifter as Eq. (6). Therefore, if Eq. (6) is additively approximated, the grouped probabilities Eq. (7) are reproduced within an additive-error through the inverse Fourier transform [40].